# Bayesian Experimental Design for Model Discrepancy Calibration: An Auto-Differentiable Ensemble Kalman Inversion Approach

### v1

@@ 1 Introduction
Bayesian experimental design (BED) provides a principled framework for optimizing data acquisition by leveraging Bayesian inference to guide the design process [1, 2, 3]. Rooted in information theory, BED aims to identify experimental conditions that maximize the information gain about target latent variables, typically by minimizing the expected posterior entropy or maximizing the expected Kullback–Leibler divergence between the prior and posterior distributions [4, 5, 6]. This probabilistic approach not only ensures that experimental designs are tailored to the specific goals of the study but also allows for adaptive, sequential experimentation, where newly acquired data refine subsequent designs [7, 8, 9, 10]. The flexibility and efficiency of BED have led to its successful application across diverse fields, including psychology [11, 12], bioinformatics [13, 14], active learning [15, 16], medical imaging [17], and engineering systems [18, 19, 20, 20, 21].
@@
However, the practical implementation of BED is often hindered by model discrepancy, a phenomenon arising from the mismatch between the predictive model and the true physical system [22]. This discrepancy introduces systematic errors in the likelihood calculation, undermining the reliability of the posterior distribution and leading to suboptimal experimental designs. Specifically, model discrepancy can result in: (i) biased parameter estimates, as the inference process is based on an incorrect likelihood; and (ii) repeated selection of similar designs, yielding datasets that lack diversity and fail to capture the true system dynamics, ultimately diminishing the effectiveness of the entire experimental design process [23, 24]. One prominent approach involves introducing correction terms and corresponding error parameters to approximate model discrepancy, thereby improving the fidelity of the likelihood function used in BED [25].
@@
Recent progresses in data-driven modeling techniques have significantly advanced the calibration of model discrepancy by leveraging some empirically chosen data to refine model predictions [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Although a BED framework has the potential of systematically identifying informative data for the calibration of some unknown parameters, the model discrepancy is often of a strongly nonlinear form and characterized by high-dimensional unknown parameters [39, 40], which poses a key challenge to standard Bayesian inference methods. While recent work has demonstrated that designs and data optimized for low-dimensional physical parameters of interest can also inform the calibration of the model discrepancy, additional evaluations with an empirically chosen threshold are often required to exclude misleading data and ensure robust performance [16]. On the other hand, classical Bayesian inference methods for information gain estimation consist of Monte Carlo-based approaches [41, 42, 43], such as Rao-Blackwellized Monte Carlo methods [44, 45] and nested Monte Carlo estimators [46], which are computationally expensive for high-dimensional problems and thus motivate the development of alternative and more affordable approaches, such as variational approaches [47, 48] and Laplace approximation [49, 50].
@@
Monte Carlo (MC) methods, despite their theoretical robustness, suffer from high computational costs due to the need for nested sampling in expected information gain (EIG) estimation, leading to an exponential increase in complexity [51, 52, 53]. The significant computational cost of MC methods also imposes a challenge to the optimization methods for search good designs. Although recent stochastic gradient descent (SGD) has improved computational efficiency by estimating EIG gradients through a small number of samples [54, 55, 56], an accurate evaluation of utility function for high-dimensional unknown parameters still demands a huge amount of samples, while the approximation via a small number of random samples could still lead to non-negligible biases. In addition, methods used to overcome the instability of gradient-based optimization for optimal design under the MC framework typically suffer from high variance, increased per-sample computational costs, or scalability issues in high-dimensional settings [57, 58, 59]. On the contrary, gradient-free approaches, such as Bayesian optimization [48, 60] and evolutionary algorithms [61, 62], can be employed for design selection but generally require an excessive number of function evaluations, making them impractical for complex and large models in many real-world applications.
@@
Variational inference (VI) provides a computationally efficient alternative to MC approaches by approximating the posterior distribution through optimization [63, 64]. Such methods typically exploit reparameterization tricks that facilitate low-variance, gradient-based optimization via automatic differentiation [65, 66]. However, their effectiveness critically depends on the forward model being differentiable with respect to the parameters of interest [67, 68, 69], while gradient-free optimization methods tend to exhibit high variance, leading to slower convergence and decreased stability, particularly in high-dimensional scenarios [70, 71, 72]. More importantly, since VI is formulated as an optimization problem, extending the framework to further optimize design variables results in a nested optimization structure. In such cases, the inner VI optimization can hinder gradient propagation to the outer level optimization, while also introducing additional computational burden and numerical instability. To enhance computational efficiency in the nested design optimization, recent approaches have focused on frameworks that bypass direct differentiation with respect to design variables by jointly optimizing the posterior approximation and design variables through maximizing a lower bound on the information gain, thereby avoiding direct divergence computations and achieving more stable gradient estimates [55]. However, like all variational inference methods, the effectiveness of these strategies depends on the quality of the variational approximation, necessitating careful selection of the approximating family [73, 74].
@@
Laplace approximation also serves as a computationally efficient alternative to MC methods and has been explored in the Bayesian optimal design. The core idea of the Laplace approximation is to approximate the posterior distribution by a Gaussian distribution centered around the maximum a posteriori (MAP) estimate [50]. This Gaussian assumption allows one to replace the otherwise intractable integration with a local second-order Taylor expansion in closed-form, which transforms the inner-loop Bayesian update, originally formulated as an integration problem, into an optimization problem [75]. This change can significantly reduce the computational cost of posterior estimation, thereby lowering the per-iteration expense of the outer-loop design optimization [76, 77]. However, the effectiveness of the Laplace approximation relies on the local Gaussianity assumption, which may break down for multimodal or highly skewed posteriors. Recent developments in Integrated Nested Laplace Approximation (INLA, [78, 79]) have been shown to mitigate these issues by providing more accurate posterior estimates, even in complex models. However, by transforming the inner posterior update into an optimization problem, these methods inherently introduce a nested (or bilevel) optimization structure, similar to variational approaches. This nested structure poses additional challenges for the optimization algorithms, as the outer-loop design optimization must now account for the computational cost and potential convergence issues of the inner-loop optimization.
@@
Instead of employing variational inference or Laplace approximation, we aim to provide a new perspective for BED, by exploring ensemble Kalman inversion (EKI, [80]) to build a framework that efficiently identifies optimal designs and data for the calibration of model discrepancy, which is often parametrized by high-dimensional unknown parameters. Unlike the VI or Laplace approximation methods that rely on the gradient descent optimization, ensemble Kalman inversion is derivative-free, parallelizable, and robust with noises of the data and chaos or uncertainties of the system [81, 82]. In addition, empirical results [83] also demonstrated that EKI is efficient and robust to handle high-dimensional and strongly nonlinear problems, despite its Gaussian assumption. In the past decade, developments have been made to enhance Kalman inversion methods in establishing the theoretical foundation [84, 85, 86]. In the context of experimental design, it is worth noting that if the inner-loop Bayesian update is fully differentiable, then the outer-loop design optimization can leverage efficient gradient-based methods without being hindered by the challenges typically associated with nested optimization. Recent work on auto-differentiable Ensemble Kalman Filters (AD-EnKF, [69]) pioneered the development of auto-differentiable ensemble Kalman methods and inspired us to explore an auto-differentiable EKI for the inner optimization problem of BED. In fact, for any other bilevel optimization problems, employing EKI in the inner optimization loop can facilitate adopting existing tools of differentiable programming for the outer-loop optimization and potentially offer computational benefits and more robust optimization performance.
@@ 2 Methodology
Bayesian experimental design (BED) is a well-established approach that relies on computational models to optimally design experiments and systematically identify informative data. In practice, model discrepancy is inevitable in many real-world applications, which accounts for one of the key challenges in BED. Although the recent developments of residual learning [37] are promising for characterizing the model discrepancy of BED, they often involve neural networks with a large number of unknown coefficients, for which MC methods via a full Bayesian perspective are often computationally infeasible. Unlike MC methods, variational methods provide more efficient, gradient-based alternatives for approximating the posterior distribution. However, when further optimizing the design based on information entropy in the out-loop, the direct application of gradient-based variational methods in the inner-loop remains challenging, limiting their efficiency in practical BED implementations. On the other hand, Laplace approximation [50, 79, 87, 76, 88] has been explored to approximately solve the Bayesian inference problem involved in BED, while the Gaussian assumption could prevent its success in addressing the neural-network-based model discrepancy that often implies strong nonlinearity.
The general goal of this work is to leverage an ensemble Kalman method that is entirely gradient-free, enabling efficient posterior updating for high-computational model discrepancy parameters in Bayesian experimental design. It is much more computationally efficient than the MC methods. Compared to variational inference methods, it only demands an ensemble of forward model evaluations and matrix/vector manipulations, which facilitates an implementation that supports automatic differentiation and allows efficient tracking of the information gain gradient with respect to the experimental design. It is worth noting that ensemble Kalman methods have been empirically shown capable of handling strong nonlinear problems in the data assimilation community, despite the Gaussian assumption behind the Kalman filter.
@@ 2.1 Bayesian Experimental Design
The Bayesian experimental design [3, 9, 89] provides a general framework to systematically seek the optimal design by solving the optimization problem:
@@
where $p(\bm{\theta}|\mathbf{y},\mathbf{d})$ is the posterior distribution of $\bm{\theta}$ given a design $\mathbf{d}$ and the data $\mathbf{y}$. Note that data $\mathbf{y}$ could be either from the actual experimental measurement or the numerical model simulation. The use of KL divergence as the utility function leads to the definition of expected information gain (EIG, [2]) that integrates the information gain over all possible predicted data:
@@ 2.2 Ensemble-based Utility Function
With the initial ensemble $\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1}$ and the updated ensemble $\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}$ after $K$ times of EKI iterations, we can approximate the KL divergence $D_{\textrm{KL}}^{e}$ between the posterior and prior distributions of $\bm{\theta}$ by treating both ensembles as multivariate Gaussian:
$\begin{gathered}D_{\textrm{KL}}^{e}(p(\bm{\theta}|\mathbf{y})||p(\bm{\theta}))\approx\tilde{D}_{\textrm{KL}}(\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}||\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1})=\\ \frac{1}{2}\left[\text{tr}\left(\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\Sigma^{\bm{\theta}\bm{\theta}}_{K}\right)-d_{\bm{\theta}}+\ln\left(\frac{\det\Sigma^{\bm{\theta}\bm{\theta}}_{0}}{\det\Sigma^{\bm{\theta}\bm{\theta}}_{K}}\right)\right.\left.+\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)^{\top}\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)\right].\end{gathered}$ (7)
@@
In summary, we employ EKI to update the ensemble of unknown parameters $\bm{\theta}$, transforming the initial ensemble $\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1}$ into the updated ones $\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}$, informed by the observation data $\mathbf{y}$ and model prediction $\mathbf{g}^{(j)}_{n}$. The divergence between these two sets of ensembles provides an approximation of the information gain $D_{\textrm{KL}}^{e}$ for the given observation data $\mathbf{y}$. Within the context of experimental design, the design variable $\mathbf{d}$ dictates the observations $\mathbf{y}(\mathbf{d})$ and the model predictions $\mathbf{g}^{(j)}_{n}(\bm{\theta}_{n}^{(j)},\mathbf{d})$, thereby establishing a mapping $\mathbf{d}\mapsto D_{\text{KL}}^{e}$. The ensemble-based approximation of KL divergence in Eq. (7) provides an efficient means to quantify the informativeness of the data $\mathbf{y}$ from a design $\mathbf{d}$. This formulation of KL divergence can be directly integrated into the EIG, and its evaluation can be efficiently implemented with parallel computing.

### v2

@@ 1 Introduction
Bayesian experimental design (BED) provides a principled framework for optimizing data acquisition by leveraging Bayesian inference to guide the design process [chaloner_bayesian_1995, lindley_measure_1956, lindley_bayesian_1972]. Rooted in information theory, BED aims to identify experimental conditions that maximize the information gain about target latent variables, typically by minimizing the expected posterior entropy or maximizing the expected Kullback–Leibler divergence between the prior and posterior distributions [raiffa_applied_2000, degroot_optimal_2005, sebastiani_maximum_2000]. This probabilistic approach not only ensures that experimental designs are tailored to the specific goals of the study but also allows for adaptive, sequential experimentation, where newly acquired data refine subsequent designs [ryan_review_2016, jones_bayes_2016, rainforth_modern_2023, huan_optimal_2024]. The flexibility and efficiency of BED have led to its successful application across diverse fields, including psychology [myung_tutorial_2013, westfall_statistical_2014], bioinformatics [flynn_bayesian_2010, thompson_integrating_2023], active learning [golovin_near-optimal_2010, yang_active_2025], medical imaging [go2025sequential], and engineering systems [papadimitriou_optimal_2004, arroyave_perspective_2022, wu2023fast, chu2025bayesian].
@@
However, the practical implementation of BED is often hindered by model discrepancy, a phenomenon arising from the mismatch between the predictive model and the true physical system [kennedy_bayesian_2001]. This discrepancy introduces systematic errors in the likelihood calculation, undermining the reliability of the posterior distribution and leading to suboptimal experimental designs. Specifically, model discrepancy can result in: (i) biased parameter estimates, as the inference process is based on an incorrect likelihood; and (ii) repeated selection of similar designs, yielding datasets that lack diversity and fail to capture the true system dynamics, ultimately diminishing the effectiveness of the entire experimental design process [grunwald_inconsistency_2017, catanach_metrics_2023]. One prominent approach involves introducing correction terms and corresponding error parameters to approximate model discrepancy, thereby improving the fidelity of the likelihood function used in BED [feng_optimal_2015].
@@
Recent progresses in data-driven modeling techniques have significantly advanced the calibration of model discrepancy by leveraging some empirically chosen data to refine model predictions [kutz2016dynamic, peherstorfer2016data, wang2017physics, wu2018physics, duraisamy2019turbulence, montans_data-driven_2019, raissi2019physics, brunton2020machine, li2020fourier, lu2021learning, kochkov2021machine, levine2022framework, chen2025neural]. Although a BED framework has the potential of systematically identifying informative data for the calibration of some unknown parameters, the model discrepancy is often of a strongly nonlinear form and characterized by high-dimensional unknown parameters [chen2024cgnsde, dong2024data], which poses a key challenge to standard Bayesian inference methods. While recent work has demonstrated that designs and data optimized for low-dimensional physical parameters of interest can also inform the calibration of the model discrepancy, additional evaluations with an empirically chosen threshold are often required to exclude misleading data and ensure robust performance [yang_active_2025]. On the other hand, classical Bayesian inference methods for information gain estimation consist of Monte Carlo-based approaches [metropolis_equation_1953, gelfand_sampling-based_1990, hastings_monte_1970], such as Rao-Blackwellized Monte Carlo methods [gal_deep_2017, foster_variational_2021] and nested Monte Carlo estimators [rainforth_nesting_2018], which are computationally expensive for high-dimensional problems and thus motivate the development of alternative and more affordable approaches, such as variational approaches [dong_variational_2025, foster_variational_2019] and Laplace approximation [laplace_memoir_1986, tierney_accurate_1986].
@@
Monte Carlo (MC) methods, despite their theoretical robustness, suffer from high computational costs due to the need for nested sampling in expected information gain (EIG) estimation, leading to an exponential increase in complexity [wakefield_expected_1994, palmer_bayesian_1998, robert_monte_2004]. The significant computational cost of MC methods also imposes a challenge to the optimization methods for search good designs. Although recent stochastic gradient descent (SGD) has improved computational efficiency by estimating EIG gradients through a small number of samples [huan_gradient-based_2014, foster_unified_2020, carlon_nesterov-aided_2020], an accurate evaluation of utility function for high-dimensional unknown parameters still demands a huge amount of samples, while the approximation via a small number of random samples could still lead to non-negligible biases. In addition, methods used to overcome the instability of gradient-based optimization for optimal design under the MC framework typically suffer from high variance, increased per-sample computational costs, or scalability issues in high-dimensional settings [giles_multilevel_2008, snyder_obstacles_2008, goda_unbiased_2022]. On the contrary, gradient-free approaches, such as Bayesian optimization [foster_variational_2019, kleinegesse_efficient_2019] and evolutionary algorithms [hamada_finding_2001, price_induced_2018], can be employed for design selection but generally require an excessive number of function evaluations, making them impractical for complex and large models in many real-world applications.
@@
Variational inference (VI) provides a computationally efficient alternative to MC approaches by approximating the posterior distribution through optimization [blei_variational_2017, ormerod_explaining_2010]. Such methods typically exploit reparameterization tricks that facilitate low-variance, gradient-based optimization via automatic differentiation [rezende_stochastic_nodate, chen_auto-encoding_2024]. However, their effectiveness critically depends on the forward model being differentiable with respect to the parameters of interest [krishnan_structured_2017, poole_variational_2019, chen_autodifferentiable_2022], while gradient-free optimization methods tend to exhibit high variance, leading to slower convergence and decreased stability, particularly in high-dimensional scenarios [williams_simple_1992, ranganath_black_2013, tucker_rebar_2017]. More importantly, since VI is formulated as an optimization problem, extending the framework to further optimize design variables results in a nested optimization structure. In such cases, the inner VI optimization can hinder gradient propagation to the outer level optimization, while also introducing additional computational burden and numerical instability. To enhance computational efficiency in the nested design optimization, recent approaches have focused on frameworks that bypass direct differentiation with respect to design variables by jointly optimizing the posterior approximation and design variables through maximizing a lower bound on the information gain, thereby avoiding direct divergence computations and achieving more stable gradient estimates [foster_unified_2020]. However, like all variational inference methods, the effectiveness of these strategies depends on the quality of the variational approximation, necessitating careful selection of the approximating family [sobolev_importance_2019, ambrogioni_automatic_2021].
@@
Laplace approximation also serves as a computationally efficient alternative to MC methods and has been explored in the Bayesian optimal design. The core idea of the Laplace approximation is to approximate the posterior distribution by a Gaussian distribution centered around the maximum a posteriori (MAP) estimate [tierney_accurate_1986]. This Gaussian assumption allows one to replace the otherwise intractable integration with a local second-order Taylor expansion in closed-form, which transforms the inner-loop Bayesian update, originally formulated as an integration problem, into an optimization problem [dodds_robust_2005]. This change can significantly reduce the computational cost of posterior estimation, thereby lowering the per-iteration expense of the outer-loop design optimization [beck_fast_2018, ryan_fully_2015]. However, the effectiveness of the Laplace approximation relies on the local Gaussianity assumption, which may break down for multimodal or highly skewed posteriors. Recent developments in Integrated Nested Laplace Approximation (INLA, [rue_approximate_2009, long_fast_2013]) have been shown to mitigate these issues by providing more accurate posterior estimates, even in complex models. However, by transforming the inner posterior update into an optimization problem, these methods inherently introduce a nested (or bilevel) optimization structure, similar to variational approaches. This nested structure poses additional challenges for the optimization algorithms, as the outer-loop design optimization must now account for the computational cost and potential convergence issues of the inner-loop optimization.
@@
Instead of employing variational inference or Laplace approximation, we aim to provide a new perspective for BED, by exploring ensemble Kalman inversion (EKI, [iglesias_ensemble_2013]) to build a framework that efficiently identifies optimal designs and data for the calibration of model discrepancy, which is often parametrized by high-dimensional unknown parameters. Unlike the VI or Laplace approximation methods that rely on the gradient descent optimization, ensemble Kalman inversion is derivative-free, parallelizable, and robust with noises of the data and chaos or uncertainties of the system [schneider2021learning, wu2024learning]. In addition, empirical results [kovachki_ensemble_2019] also demonstrated that EKI is efficient and robust to handle high-dimensional and strongly nonlinear problems, despite its Gaussian assumption. In the past decade, developments have been made to enhance Kalman inversion methods in establishing the theoretical foundation [schillings2017analysis, ding2021ensemble, calvello2022ensemble]. In the context of experimental design, it is worth noting that if the inner-loop Bayesian update is fully differentiable, then the outer-loop design optimization can leverage efficient gradient-based methods without being hindered by the challenges typically associated with nested optimization. Recent work on auto-differentiable Ensemble Kalman Filters (AD-EnKF, [chen_autodifferentiable_2022]) pioneered the development of auto-differentiable ensemble Kalman methods and inspired us to explore an auto-differentiable EKI for the inner optimization problem of BED. In fact, for any other bilevel optimization problems, employing EKI in the inner optimization loop can facilitate adopting existing tools of differentiable programming for the outer-loop optimization and potentially offer computational benefits and more robust optimization performance.
@@ 2 Methodology
The difference between the true system and a physics-based model, known as model discrepancy, is ubiquitous in the modeling and simulation of complex dynamical systems [kennedy_bayesian_2001, brynjarsdottir2014learning]. Various data-driven methods have been explored to learn model discrepancy, improving state forecasts or parameter estimates [luo2020accounting, dong2025stochastic]. Existing data-driven methods on the calibration of model discrepancy often rely on empirical choices of the training data rather than systematically designing it. In contrast to the physical parameters of a model, which typically reside in a low-dimensional vector space, the model discrepancy is an element of an infinite-dimensional function space [wu2024learning]. Consequently, approximating this function with methods such as neural networks leads to a high-dimensional coefficient space, which poses a significant challenge for systematically identifying optimal training data using standard BED methods [yang_active_2025]. In this work, we aim to build on the AD-EKI and develop an efficient framework to identify optimal designs that guide active data acquisition for the calibration of the model discrepancy.
The true system in a general form can be written as:
The neural network NN is trained using measurement data $\mathbf{y}$, which can be viewed as a function of the experimental design $\mathbf{d}$ in the context of BED. In this work, the design $\mathbf{d}$ corresponds to the spatiotemporal coordinate $(\mathbf{z},t)$. The measurement $\mathbf{y}$ is the observation of variable $\mathbf{u}$ that represents the state of the physical system at the design $\mathbf{d}$, i.e., $\mathbf{y}=\mathbf{u}(\mathbf{d})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\epsilon}})$ represents the measurement noise which is assumed to be Gaussian in this work when observing the state from the true system.
@@ 2.1 Bayesian Experimental Design
The Bayesian experimental design [lindley_bayesian_1972, rainforth_modern_2023, huan_optimal_2024] provides a general framework to systematically seek the optimal design by solving the optimization problem:
@@
where $p(\bm{\theta}|\mathbf{y},\mathbf{d})$ is the posterior distribution of $\bm{\theta}$ given a design $\mathbf{d}$ and the data $\mathbf{y}$. Note that data $\mathbf{y}$ could be either from the actual experimental measurement or the numerical model simulation. The use of KL divergence as the utility function leads to the definition of expected information gain (EIG, [lindley_measure_1956]) that integrates the information gain over all possible predicted data:
@@
The performance of BED relies on the accuracy of forward model [rainforth_modern_2023, catanach_metrics_2023], while the model discrepancy [kennedy_bayesian_2001, feng_optimal_2015] is inevitable in most real-world applications. Although the recent developments of residual learning [levine2022framework, yang_active_2025] are promising for characterizing the model discrepancy of BED, they often involve neural networks with a large number of unknown coefficients, for which systematically gathering informative training data via a full Bayesian perspective (e.g., MC methods) is often computationally infeasible. Unlike MC methods, variational methods provide more efficient, gradient-based alternatives for approximating the posterior distribution. However, when further optimizing the design based on information entropy in the outer-loop, the direct application of gradient-based variational methods in the inner-loop remains challenging, limiting their efficiency in practical BED implementations and potentially leading to an inconsistent formulation compared to the original bilevel optimization problem. On the other hand, Laplace approximation [tierney_accurate_1986, long_fast_2013, long_laplace_2015, beck_fast_2018, bartuska2025laplace] has been explored to approximately solve the Bayesian inference problem involved in BED, while the intrinsic Gaussian assumption could be too restrictive in addressing the neural-network-based model discrepancy that often implies strong nonlinearity. Therefore, we aim to propose an alternative approach to the bilevel optimization problem of BED for model discrepancy, by developing an ensemble Kalman method for the inner-loop optimization that can handle nonlinearity and supports automatic differentiation, to facilitate computing gradients of information gain with respect to the experimental design for the outer-loop optimization.
@@ 2.2 Ensemble-based Utility Function
With the initial ensemble $\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1}$ and the updated ensemble $\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}$ after $K$ times of EKI iterations, we follow the standard information-based BED paradigm [chaloner_bayesian_1995, lindley_bayesian_1972] and quantify the information gain using the KL divergence between these two ensembles. By assuming each ensemble is drawn from a multivariate Gaussian distribution, this ensemble-based KL admits a closed-form approximation:
$\begin{gathered}D_{\textrm{KL}}(p(\bm{\theta}|\mathbf{y})||p(\bm{\theta}))\approx\tilde{D}_{\textrm{KL}}(\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}||\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1})=\\ \frac{1}{2}\left[\text{tr}\left(\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\Sigma^{\bm{\theta}\bm{\theta}}_{K}\right)-d_{\bm{\theta}}+\ln\left(\frac{\det\Sigma^{\bm{\theta}\bm{\theta}}_{0}}{\det\Sigma^{\bm{\theta}\bm{\theta}}_{K}}\right)\right.\left.+\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)^{\top}\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)\right].\end{gathered}$ (9)
@@
In summary, we employ EKI to update the ensemble of unknown parameters $\bm{\theta}$, transforming the initial ensemble $\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1}$ into the updated ones $\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}$, informed by the observation data $\mathbf{y}$ and model prediction $\mathbf{g}^{(j)}_{n}$. The divergence between these two sets of ensembles provides an approximation $\tilde{D}_{\text{KL}}$ of the information gain $D_{\textrm{KL}}$ for the given observation data $\mathbf{y}$. Within the context of experimental design, the design variable $\mathbf{d}$ dictates the observations $\mathbf{y}(\mathbf{d})$ and the model predictions $\mathbf{g}^{(j)}_{n}(\bm{\theta}_{n}^{(j)},\mathbf{d})$, thereby establishing a mapping $\mathbf{d}\mapsto\tilde{D}_{\text{KL}}$. The ensemble-based approximation of KL divergence in Eq. (9) provides an efficient means to quantify the informativeness of the data $\mathbf{y}$ from a design $\mathbf{d}$. This formulation of KL divergence can be directly integrated into the EIG, and its evaluation can be efficiently implemented with parallel computing.